### 01 Feb 2024 v1

@@ I Introduction
In 2012, the 125 GeV Higgs boson discovered by the Large Hadron Collider (LHC) mh-ATLAS ; mh-CMS , the measured mass of the Higgs boson now is PDG1 $m_{h}=125.25\pm 0.17\>{\rm{GeV}}.$ The discovery of the Higgs boson marked a huge success for the Standard Model (SM), but it did not stop the search for New Physics (NP) at the LHC, and one of them was the search for lighter scalar particles.
@@
As one of the extensions of the SM, the $\mu$ from $\nu$ supersymmetric standard model ($\mu$$\nu$SSM) mnSSM ; mnSSM1 ; mnSSM1-1 ; parameter-space ; mnSSM2 ; mnSSM2-1 ; Zhang1 ; Zhang2 can solve the $\mu$ problem m-problem of the minimal supersymmetric standard model (MSSM) MSSM ; MSSM1 ; MSSM2 ; MSSM3 ; MSSM4 , through introducing three singlet right-handed neutrino superfields $\hat{\nu}_{i}^{c}$ ($i=1,2,3$). The neutrino superfields lead the mixing of the neutral components of the Higgs doublets with the right-handed sneutrinos, that is different from the Higgs sector of the MSSM. The mixing can change the Higgs couplings, and influence the decay processes of the Higgs bosons. In addition, we also introduce CP-violation, and we also get a lighter Higgs at $\sim$ 95 GeV with a suitable parameter space.
The paper is organized as follows. In Sec. II, we introduce the CPV $\mu\nu$SSM briefly, about the superpotential and the CPV phases. In Sec. III, we study the excess at 95 GeV in the CPV $\mu\nu$SSM. Sec. IV and Sec. V, we show the numerical analysis and the conclusion.
@@ II The Model
The superpotential of the $\mu\nu$SSM contains Yukawa couplings for neutrinos, two additional types of terms involving the Higgs doublet superfields $\hat{H}_{u}$ and $\hat{H}_{d}$, and the right-handed neutrino superfields $\hat{\nu}_{i}^{c}$mnSSM
@@
where $\hat{H}_{u}^{T}=\Big{(}{\hat{H}_{u}^{+},\hat{H}_{u}^{0}}\Big{)}$, $\hat{H}_{d}^{T}=\Big{(}{\hat{H}_{d}^{0},\hat{H}_{d}^{-}}\Big{)}$, $\hat{Q}_{i}^{T}=\Big{(}{{{\hat{u}}_{i}},{{\hat{d}}_{i}}}\Big{)}$, $\hat{L}_{i}^{T}=\Big{(}{{{\hat{\nu}}_{i}},{{\hat{e}}_{i}}}\Big{)}$ (the index $T$ denotes the transposition) represent the MSSM-like doublet Higgs superfields, and $\hat{u}_{i}^{c}$, $\hat{d}_{i}^{c}$, and $\hat{e}_{i}^{c}$ are the singlet up-type quark, down-type quark and charged lepton superfields, respectively. In addition, $Y_{u,d,e,\nu}$, $\lambda$, and $\kappa$ are dimensionless matrices, a vector, and a totally symmetric tensor. $a,b=1,2$ are SU(2) indices with antisymmetric tensor $\epsilon_{12}=1$, and $i,j,k=1,2,3$ are generation indices. The CP is violated by the parameter $\lambda_{i}$, the CP-violating phase is $\phi_{\lambda_{i}}$.
In the superpotential, if the scalar potential is such that nonzero vacuum expectative values (VEVs) of the scalar components ($\tilde{\nu}_{i}^{c}$) of the singlet neutrino superfields $\hat{\nu}_{i}^{c}$ are induced, the effective bilinear terms $\epsilon_{ab}\varepsilon_{i}\hat{H}_{u}^{b}\hat{L}_{i}^{a}$ and $\epsilon_{ab}\mu\hat{H}_{d}^{a}\hat{H}_{u}^{b}$ are generated, with $\varepsilon_{i}=Y_{\nu_{ij}}\left\langle{\tilde{\nu}_{j}^{c}}\right\rangle$ and $\mu={\lambda_{i}}\left\langle{\tilde{\nu}_{i}^{c}}\right\rangle$, once the electroweak symmetry is broken. The last term Eq. (1) generates the effective Majorana masses for neutrinos at the electroweak scale. Therefore, the $\mu\nu$SSM can generate three tiny neutrino masses at the tree level through TeV scale seesaw mechanism mnSSM ; neutrino-mass ; neu-mass1 ; neu-mass2 ; neu-mass3 ; neu-mass4 ; neu-mass5 ; neu-mass6 .
It is worth explaining why TeV scale seesaw was chosen. Through a seesaw on the scale of the Grand Unified Theory (GUT), one can get the Yukawa couplings of order one for neutrinos. But we know that the Yukawa coupling of the electron is on the order of $10^{-6}$, and the Yukawa couplings of neutrinos can also be around on the order of $10^{-6}$ instead of one. In the TeV scale seesaw, this is sufficient to reproduce the neutrino mass, if the Yukawa coupling of the neutrino is of the same order as the Yukawa coupling of the electron mnSSM . Here it is important to note that the VEVs of the left-handed sneutrinos $\upsilon_{\nu_{i}}$ are generally small. We know that the Dirac masses for the neutrinos $m_{D_{i}}=Y_{\upsilon_{\nu_{i}}}\upsilon_{u}\lesssim 10^{-4}$ GeV in the TeV scale seesaw. So we can get an estimate of the VEVs of the left-handed sneutrinos, $\upsilon_{\nu_{i}}\lesssim m_{D_{i}}\lesssim 10^{-4}$ GeV, which means that $\upsilon_{\nu_{i}}\ll\upsilon_{d},\upsilon_{u}$ mnSSM ; mnSSM1 .
The general soft SUSY-breaking terms of the $\mu\nu$SSM are given by
@@
$H_{d}^{0}=e^{i\phi_{\upsilon_{d}}}(\frac{h_{d}+iP_{d}}{\sqrt{2}}+\upsilon_{d}),\qquad\;\tilde{\nu}_{i}=\frac{(\tilde{\nu}_{i})^{\Re}+i(\tilde{\nu}_{i})^{\Im}}{\sqrt{2}}+\upsilon_{\nu_{i}},$
$H_{u}^{0}=e^{i\phi_{\upsilon_{u}}}(\frac{h_{u}+iP_{u}}{\sqrt{2}}+\upsilon_{u}),\qquad\tilde{\nu}_{i}^{c}=e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}(\frac{(\tilde{\nu}_{i}^{c})^{\Re}+i(\tilde{\nu}_{i}^{c})^{\Im}}{\sqrt{2}}+\upsilon_{\nu_{i}^{c}}),$ (4)
@@
In supersymmetric (SUSY) extensions of the SM, the R-parity of a particle is defined as $R=(-1)^{L+3B+2S}$ MSSM ; MSSM1 ; MSSM2 ; MSSM3 ; MSSM4 . R-parity is violated if either the baryon number ($B$) or lepton number ($L$) is not conserved, where $S$ denotes the spin of concerned component field. The last two terms in Eq. (1) explicitly violate lepton number and R-parity. For example, if one assigns $L=1$ to the right-handed neutrino superfields, then the last term ${1\over 3}\kappa_{ijk}\hat{\nu}_{i}^{c}\hat{\nu}_{j}^{c}\hat{\nu}_{k}^{c}$ of Eq. (2) violates the lepton number by three units contrary to the $\hat{L}_{i}\hat{L}_{j}\hat{e}^{c}_{k}$ term of the R-parity violating MSSM which shows $\Delta L=1$ effect. R-parity breaking implies that the lightest supersymmetric particle (LSP) is no longer stable.
II.1 The $\mu\nu$SSM Higgs potential
The neutral scalar potential of the tree-level can be written as:
@@ II.1 The $\mu\nu$SSM Higgs potential
where $G^{2}=g_{1}^{2}+g_{2}^{2}$. The tree-level neutral scalar potential include the usual soft terms and $D-$ and $F-$term contributions. In this work, we take all parameters in the potential ara real.
By using the effective potential methods epm ; epm1 ; epm2 ; epm3 ; epm4 ; epm5 ; epm6 ; epm7 ; epm8 ; epm9 ; epm10 ; epm11 ; epm12 ; epm13 ; epm14 ; epm15 ; epm16 , one can get the one-loop efffective potential:
@@
Considering the one-loop effective potential, the Higgs potential can be writtrn as:
@@
we can calculte the minimization conditions of the potential and the Higgs masses in this work.
@@ II.2 Higgs masses and CPV phases
In the $\mu\nu$SSM, the left- & right-handed sneutrino VEVs lead to the mixing of the neutral components of the Higgs doublets with the left- &\&& right-handed sneutrinos producing an $8\times 8$ CP-even neutral scalar mass matrix, which can be seen in Refs. mnSSM1 ; mnSSM1-1 ; Zhang1 ; MASS . The mixing gives a rich phenomenology in the Higgs sector of the $\mu\nu$SSM mnSSM1 ; mnSSM2 ; mnSSM1-1 ; parameter-space ; mnSSM2-1 ; HZrr ; muon ; MASS ; HZr .
Here, we note that the Higgs doublets and right-handed sneutrinos are basically decoupled from the left-handed sneutrinos MASS , so we didn’t considering the left-handed sneutrinos in the CP-even and CP-odd scalars part.
@@
with $M_{S}^{2}$ denote the CP-even neutral scalars, $M_{P}^{2}$ is the CP-odd neutral scalars, $M_{SP}^{2}$ represents the mass submatrix for the mixing of CP-even neutral scalars and CP-odd neutral scalars.
The mass sqared matrix $M_{h}^{2}$ can be diagonalized as:
@@
We consider the radiative correcttions in mass submatrix $M_{H}^{2}$, the radiative corrections from the third fermions ($f=t,b,\tau$) and their superpartners include the two-loop leading-log effects leading-log ; leading-log1 ; leading-log2 . The CP-even neutral scalars is given as:
@@
In detail, the mass submatrix $M_{H}^{2}$ is defind by
@@
The dominating contributions of radiative corrections $\Delta_{11},\Delta_{12}$ and $\Delta_{22}$ comes from the fermions ($t,b$) and their superpartners
@@
We didn’t consider the terms containing coupling $Y_{\nu_{i}}$ and $\upsilon_{\nu_{i}}$, because these terms are very small. The radiative corrections from the top quark is give by HZrr ; radiative corr ; radiative corr1 ; radiative corr2 ; radiative corr3 ; radiative corr4 ; radiative corr5
$\Delta_{11}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}{\beta}}}{{(\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu-|\mu|^{2}\cot\beta)^{2}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})$ (23)
$\Delta_{12}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}{\beta}}}{{|\mu|^{2}\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (24)
×$\times\left(\ln{{m_{\tilde{t}_{1}^{2}}^{2}}\over{m_{\tilde{t}_{2}}^{2}}}+{{\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\right)$
$\Delta_{22}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}{\beta}}}\Bigg{\{}\ln\frac{m_{\tilde{t}_{1}}^{2}m_{\tilde{t}_{2}}^{2}}{m_{t^{4}}}+{{2|A_{t}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cos\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}\ln{{m_{\tilde{t}_{1}^{2}}^{2}}\over{m_{\tilde{t}_{2}}^{2}}}$ (25)
$+{{(2|A_{t}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta)^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2})^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})+{{1}\over{16\pi^{2}}}\ln{{m_{\tilde{t}_{1}}^{2}m_{\tilde{t}_{2}}^{2}}\over{m_{t}^{4}}}$
@@
to save space, the mass matrix and the radiative corrections are given in the Appendix A.
In the radiative corrections, the trilinear coupling $A_{t}=|A_{t}|e^{i\phi_{A_{t}}}$ can be complex. These seven independent phase which have been defined as:
$\phi_{\lambda},\phi_{A_{t}},\phi_{\upsilon_{u}},\phi_{\upsilon_{d}},\phi_{\upsilon_{\nu_{i}^{c}}}$ (27)

### 06 Mar 2024 v2

@@ I Introduction
In 2012, the 125 GeV Higgs boson discovered by the Large Hadron Collider (LHC) mh-ATLAS ; mh-CMS ; the measured mass of the Higgs boson now is PDG1 $m_{h}=125.25\pm 0.17\>{\rm{GeV}}$. The discovery of the Higgs boson marked a huge success for the standard model (SM), but it did not stop the search for new physics (NP) at the LHC, and one of them was the search for lighter scalar particles.
@@
As one of the extensions of the SM, the $\mu$-from-$\nu$ supersymmetric standard modelmnSSM ; mnSSM1 ; mnSSM1-1 ; parameter-space ; mnSSM2 ; mnSSM2-1 ; Zhang1 ; Zhang2 can solve the $\mu$ problem m-problem of the minimal supersymmetric standard model (MSSM) MSSM ; MSSM1 ; MSSM2 ; MSSM3 ; MSSM4 , through introducing three singlet right-handed neutrino superfields $\hat{\nu}_{i}^{c}$ ($i=1,2,3$). The neutrino superfields lead the mixing of the neutral components of the Higgs doublets with the right-handed sneutrinos, that is different from the Higgs sector of the MSSM. The mixing can change the Higgs couplings and influence the decay processes of the Higgs bosons. In addition, we also introduce $CP$ violation, and we also get a lighter Higgs at $\sim$ 95 GeV with a suitable parameter space.
The paper is organized as follows. In Sec. II, we introduce the $CP$-violating $\mu$-from-$\nu$ SSM briefly, about the superpotential and the CPV phases. In Sec. III, we study the excess at 95 GeV in the CPV $\mu$-from-$\nu$ SSM. In Secs. IV and V, we show the numerical analysis and the conclusion,respectively.
@@ II The Model
The superpotential of the $\mu$-from-$\nu$ SSM contains Yukawa couplings for neutrinos, two additional types of terms involving the Higgs doublet superfields $\hat{H}_{u}$ and $\hat{H}_{d}$, and the right-handed neutrino superfields $\hat{\nu}_{i}^{c}$ mnSSM :
@@
where $\hat{H}_{u}^{T}=\Big{(}{\hat{H}_{u}^{+},\hat{H}_{u}^{0}}\Big{)}$, $\hat{H}_{d}^{T}=\Big{(}{\hat{H}_{d}^{0},\hat{H}_{d}^{-}}\Big{)}$, $\hat{Q}_{i}^{T}=\Big{(}{{{\hat{u}}_{i}},{{\hat{d}}_{i}}}\Big{)}$, and $\hat{L}_{i}^{T}=\Big{(}{{{\hat{\nu}}_{i}},{{\hat{e}}_{i}}}\Big{)}$ (the index $T$ denotes the transposition) represent the MSSM-like doublet Higgs superfields and $\hat{u}_{i}^{c}$, $\hat{d}_{i}^{c}$, and $\hat{e}_{i}^{c}$ are the singlet up-type quark, down-type quark and charged lepton superfields, respectively. In addition, $Y_{u,d,e,\nu}$, $\lambda$, and $\kappa$ are dimensionless matrices, a vector, and a totally symmetric tensor, respectively. $a,b=1,2$ are SU(2) indices with antisymmetric tensor $\epsilon_{12}=1$, and $i,j,k=1,2,3$ are generation indices. The $CP$ is violated by the parameter $\lambda_{i}$, and the $CP$-violating phase is $\phi_{\lambda_{i}}$.
In the superpotential, if the scalar potential is such that nonzero vacuum expectative values (VEVs) of the scalar components ($\tilde{\nu}_{i}^{c}$) of the singlet neutrino superfields $\hat{\nu}_{i}^{c}$ are induced, the effective bilinear terms $\epsilon_{ab}\varepsilon_{i}\hat{H}_{u}^{b}\hat{L}_{i}^{a}$ and $\epsilon_{ab}\mu\hat{H}_{d}^{a}\hat{H}_{u}^{b}$ are generated, with $\varepsilon_{i}=Y_{\nu_{ij}}\left\langle{\tilde{\nu}_{j}^{c}}\right\rangle$ and $\mu={\lambda_{i}}\left\langle{\tilde{\nu}_{i}^{c}}\right\rangle$, once the electroweak symmetry is broken. The last term in Eq. (1) generates the effective Majorana masses for neutrinos at the electroweak scale. Therefore, the $\mu$-from-$\nu$ SSM can generate three tiny neutrino masses at the tree level through TeV-scale seesaw mechanism mnSSM ; neutrino-mass ; neu-mass1 ; neu-mass2 ; neu-mass3 ; neu-mass4 ; neu-mass5 ; neu-mass6 .
It is worth explaining why the TeV-scale seesaw was chosen. Through a seesaw on the scale of the grand unified theory, one can get Yukawa couplings of the order of one for neutrinos. But we know that the Yukawa coupling of the electron is on the order of $10^{-6}$, and the Yukawa couplings of neutrinos can also be around on the order of $10^{-6}$ instead of one. In the TeV-scale seesaw, this is sufficient to reproduce the neutrino mass, if the Yukawa coupling of the neutrino is of the same order as the Yukawa coupling of the electron mnSSM . Here, it is important to note that the VEVs of the left-handed sneutrinos $\upsilon_{\nu_{i}}$ are generally small. We know that the Dirac masses for the neutrinos $m_{D_{i}}=Y_{\upsilon_{\nu_{i}}}\upsilon_{u}\lesssim 10^{-4}$ GeV in the TeV-scale seesaw. So we can get an estimate of the VEVs of the left-handed sneutrinos, $\upsilon_{\nu_{i}}\lesssim m_{D_{i}}\lesssim 10^{-4}$ GeV, which means that $\upsilon_{\nu_{i}}\ll\upsilon_{d},\upsilon_{u}$ mnSSM ; mnSSM1 .
The general soft supersymmetry-breaking terms of the $\mu$-from-$\nu$SSM are given by
@@
$H_{d}^{0}=e^{i\phi_{\upsilon_{d}}}(\frac{h_{d}+iP_{d}}{\sqrt{2}}+\upsilon_{d}),\qquad\;\tilde{\nu}_{i}=\frac{(\tilde{\nu}_{i})^{\text{Re}}+i(\tilde{\nu}_{i})^{\text{Im}}}{\sqrt{2}}+\upsilon_{\nu_{i}},$
$H_{u}^{0}=e^{i\phi_{\upsilon_{u}}}(\frac{h_{u}+iP_{u}}{\sqrt{2}}+\upsilon_{u}),\qquad\tilde{\nu}_{i}^{c}=e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}(\frac{(\tilde{\nu}_{i}^{c})^{\text{Re}}+i(\tilde{\nu}_{i}^{c})^{\text{Im}}}{\sqrt{2}}+\upsilon_{\nu_{i}^{c}}),$ (4)
@@
In supersymmetric extensions of the SM, the $R$ parity of a particle is defined as $R=(-1)^{L+3B+2S}$ MSSM ; MSSM1 ; MSSM2 ; MSSM3 ; MSSM4 . $R$ parity is violated if either the baryon number ($B$) or lepton number ($L$) is not conserved, where $S$ denotes the spin of concerned component field. The last two terms in Eq. (1) explicitly violate lepton number and $R$ parity. For example, if one assigns $L=1$ to the right-handed neutrino superfields, then the last term ${1\over 3}\kappa_{ijk}\hat{\nu}_{i}^{c}\hat{\nu}_{j}^{c}\hat{\nu}_{k}^{c}$ in Eq. (2) violates the lepton number by three units contrary to the $\hat{L}_{i}\hat{L}_{j}\hat{e}^{c}_{k}$ term of the $R$ parity violating MSSM which shows the $\Delta L=1$ effect. $R$ parity breaking implies that the lightest supersymmetric particle is no longer stable.
II.1 The $\mu$-from-$\nu$SSM Higgs potential
The neutral scalar potential of the tree level can be written as
@@ II.1 The $\mu$-from-$\nu$SSM Higgs potential
where $G^{2}=g_{1}^{2}+g_{2}^{2}$. The tree-level neutral scalar potential include the usual soft terms and $D$- and $F$-term contributions. In this work, we take all parameters in the potential are real.
By using the effective potential methods epm ; epm1 ; epm2 ; epm3 ; epm4 ; epm5 ; epm6 ; epm7 ; epm8 ; epm9 ; epm10 ; epm11 ; epm12 ; epm13 ; epm14 ; epm15 ; epm16 , one can get the one-loop effective potential:
@@
Considering the one-loop effective potential, the Higgs potential can be written as
@@
we can calculate the minimization conditions of the potential and the Higgs masses in this work.
@@ II.2 Higgs masses and CPV phases
In the $\mu$-from-$\nu$SSM, the left- and right-handed sneutrino VEVs lead to the mixing of the neutral components of the Higgs doublets with the left- and right-handed sneutrinos producing an $8\times 8$ $CP$-even neutral scalar mass matrix, which can be seen in Refs. mnSSM1 ; mnSSM1-1 ; Zhang1 ; MASS . The mixing gives a rich phenomenology in the Higgs sector of the $\mu$-from-$\nu$SSM mnSSM1 ; mnSSM2 ; mnSSM1-1 ; parameter-space ; mnSSM2-1 ; HZrr ; muon ; MASS ; HZr .
Here, we note that the Higgs doublets and right-handed sneutrinos are basically decoupled from the left-handed sneutrinos MASS , so we did not consider the left-handed sneutrinos in the $CP$-even and $CP$-odd scalars part.
@@
with $M_{S}^{2}$ denoting the $CP$-even neutral scalars, $M_{P}^{2}$ is the $CP$-odd neutral scalars, and $M_{SP}^{2}$ represents the mass submatrix for the mixing of $CP$-even neutral scalars and $CP$-odd neutral scalars.
The mass squared matrix $M_{h}^{2}$ can be diagonalized as
@@
We consider the radiative corrections in mass submatrix $M_{H}^{2}$; the radiative corrections from the third fermions ($f=t,b,\tau$) and their superpartners include the two-loop leading-log effects leading-log ; leading-log1 ; leading-log2 . The $CP$-even neutral scalars is given as:
@@
In detail, the mass submatrix $M_{H}^{2}$ is defined by
@@
The dominating contributions of radiative corrections $\Delta_{11},~{}\Delta_{12}$ and $\Delta_{22}$ comes from the fermions ($t,b$) and their superpartners
@@
We did not consider the terms containing coupling $Y_{\nu_{i}}$ and $\upsilon_{\nu_{i}}$, because these terms are very small. The radiative corrections from the top quark is given by HZrr ; radiative corr ; radiative corr1 ; radiative corr2 ; radiative corr3 ; radiative corr4 ; radiative corr5
$\Delta_{11}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}{\beta}}}{{(\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu-|\mu|^{2}\cot\beta)^{2}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})$ (23)
$\Delta_{12}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}{\beta}}}{{|\mu|^{2}\cot\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (24)
×$\times\left(\ln{{m_{\tilde{t}_{1}^{2}}^{2}}\over{m_{\tilde{t}_{2}}^{2}}}+{{\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\right)$
$\Delta_{22}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}{\beta}}}\Bigg{\{}\ln\frac{m_{\tilde{t}_{1}}^{2}m_{\tilde{t}_{2}}^{2}}{m_{t^{4}}}+{{2|A_{t}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cos\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}\ln{{m_{\tilde{t}_{1}^{2}}^{2}}\over{m_{\tilde{t}_{2}}^{2}}}$ (25)
$+{{(2|A_{t}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{\upsilon_{u}}}e^{-i\phi_{\upsilon_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta)^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2})^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})+{{1}\over{16\pi^{2}}}\ln{{m_{\tilde{t}_{1}}^{2}m_{\tilde{t}_{2}}^{2}}\over{m_{t}^{4}}}$
@@
to save space, the mass matrix and the radiative corrections are given in the Appendix.
In the radiative corrections, the trilinear coupling $A_{t}=|A_{t}|e^{i\phi_{A_{t}}}$ can be complex. These seven independent phase have been defined as
$\phi_{\lambda},\phi_{A_{t}},\phi_{\upsilon_{u}},\phi_{\upsilon_{d}},\phi_{\upsilon_{\nu_{i}^{c}}}.$ (27)